OpenStudy (anonymous):

Limit Question \[ \lim_{x\to 0}f(|x|) = \lim_{x\to0^+}f(x) \]I am wondering if this is true or not.

OpenStudy (anonymous):

I mean \(f(x) = x\), we have: \[ \lim_{x\to 0}|x| =0= \lim_{x\to 0^+}x \]But I'm wondering about more general cases.

hartnn (hartnn):

\(\lim_{x\to 0}|x| =0= \lim_{x\to 0^-}x\) also ... and for \(\lim_{x\to 0}f(|x|) \) to exist it must be \( = \lim_{x\to0^+}f(x)= \lim_{x\to0^-}f(x)\) both

OpenStudy (anonymous):

Really @hartnn ?

OpenStudy (anonymous):

I think this is true:\[ \lim_{x\to 0}\frac{1}{|x|} = \infty \]

OpenStudy (anonymous):

I'm using the function \(f(x) = 1/x\) this time.

OpenStudy (anonymous):

But doesn't: \[ \lim_{x\to 0^-}\frac{1}{x} = -\infty \]?

OpenStudy (anonymous):

Or did I mix them up?

hartnn (hartnn):

yes wio, and yes thats true even for f(x) =1/x or any f(x) if lim x->0 f(|x|) exist, then its a must that \(= \lim_{x\to0^+}f(x)= \lim_{x\to0^-}f(x) =\lim_{x\to0}f(x) \)

hartnn (hartnn):

if you are taking x->0- then |x| = -x

hartnn (hartnn):

\(\lim_{x\to 0^-}\frac{1}{-x} = \infty\)

OpenStudy (anonymous):

Why do both limits have to exist?

hartnn (hartnn):

woah! why are so many offline memebers atching this post :O :P

hartnn (hartnn):

*watching and they keep on increasing !

OpenStudy (anonymous):

Since \(|x|>0\) mean that the negative side doesn't matter?

OpenStudy (anonymous):

@hartnn Why does the negative side matter if \(|x|>0\)?

hartnn (hartnn):

my guess : some school teacher introduced OpenStudy to their class, and all of them came to check it out without logging in :P wio, f(x) = 1/x f(|x|) = 1/|x| for x->0+ , x>0 1/|x| = 1/x for x->0-, x<0 so 1/|x| = -1/x

OpenStudy (anonymous):

\[ \lim_{x\to 0}f(|x|)=L \]Okay so we have: \[ |x|<\delta \implies |f(|x|)-L)|<\epsilon \]

OpenStudy (anonymous):

@hartnn Yeah, but in this case \(f(|x|)\) exists but \(f(x)\) doesn't. So that already is against what you said.

hartnn (hartnn):

you mean at x=0 ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

You said that \(f(|x|)\) doesn't exist unless \(f(x)\) does, but that doesn't seem to be true for \(f(x)=1/x\).

hartnn (hartnn):

i thought, f(x) exist at x=0 and is = \(\infty \)

OpenStudy (anonymous):

For \[ \lim_{x\to0+} f(x)=L \]We have: \[ x<\delta \implies |f(x)-L|<\epsilon \]

OpenStudy (anonymous):

And \(0<x\) as well

OpenStudy (anonymous):

Doesn't it seem like they're equivalent?

OpenStudy (anonymous):

If you let \(y=|x|\), then they are equivalent, I think

OpenStudy (anonymous):

@hartnn Do you want to understand delta epsilon a bit better?

hartnn (hartnn):

yup, f(x) does not exist at x=0 also i meant \(= \lim_{x\to0^+}|f(x)|= \lim_{x\to0^-}|f(x)|\) i just saw i missed || in my reply i never meant " f(|x|) doesn't exist unless f(x) does, but that doesn't seem to be true for f(x)=1/x."

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @hartnn if lim x->0 f(|x|) exist, then its a must that \(= \lim_{x\to0^+}f(x)= \lim_{x\to0^-}f(x) =\lim_{x\to0}f(x) \) \(\color{blue}{\text{End of Quote}}\) This is tha part that confused me...

hartnn (hartnn):

yes, i missed || there too :( \(= \lim_{x\to0^+}|f(x)|= \lim_{x\to0^-}|f(x)| =\lim_{x\to0}|f(x)|\)

OpenStudy (anonymous):

Oh, ok. I was thinking about doing a delta epsilon tutorial

hartnn (hartnn):

go on! many users will be benefited, including me :)

OpenStudy (anonymous):

Okay, I think I'll write it up in a new question, but I probably better plan it out a bit

hartnn (hartnn):

did we conclude on your question ?

OpenStudy (anonymous):

Yeah, I'll close this.

OpenStudy (anonymous):

I'm just using this to preview mathjax